You release a 1.4-kg ball from rest at a height of 2 m above soft ground. Once it has come to rest, the ball has made an 8-cm deep hole in the ground.
What is the magnitude of the force that the ground exerts on the ball while it is stopping?
1
Expert's answer
2019-10-09T10:41:29-0400

The change of energy is equal to work done, so


"\Delta E=W"

"mgh=Fd"

Therefore, the magnitude of the force


"F=\frac{mgh}{d}"

"=\frac{1.4\:\rm kg\times 9.8\: m/s^2\times 2\: m}{0.08\:\rm m}=343\:\rm N"